OpenStudy (anonymous):

solve for x logx+log(x-3)=log(3x) is the answer x=0?

OpenStudy (zzr0ck3r):

\(\log(x)+\log(x-3)=\log(x(x-3))\) so taking both sides to the power of 10 we see \(x(x-3)=3x\) solve for x

OpenStudy (anonymous):

the options are x=0 x=0,6 x=6 or there are no solutions

OpenStudy (zzr0ck3r):

sove for x

OpenStudy (zzr0ck3r):

\(x^2-3x-3x=0\\x^2-6x=0\\x(x-6)=0\\so\\x=0 \ or \ x=6\)but x cant be 0

OpenStudy (anonymous):

\[\log_{3x}-\log_{x}=\log_{(x-3)} \] \[\log_{(3x/x)}=\log_{3}=\log_{(x-3)} \] \[x-3=3\] \[x=6\]

OpenStudy (zzr0ck3r):

\(x^2-3x-3x=0\\x^2-6x=0\\x(x-6)=0\\so\\x=0 \ or \ x=6\) but x cant be 0, x = 6

OpenStudy (anonymous):

oh I see ok I get it now thanks